Dec. 22 – Jan. 3

– Five stores provided ~65.5 hours of luminosity
– T-932 (Diamond) began taking beam at MTest
– Store 8388 quenched due to a vacuum problem
– Cryo system technicians began a warm up of B1 to change out a spool piece
– NuMI and MiniBooNE continued to take beam as it was available
– Beam to T-932 stopped due to accesses into the transfer hall
– Machines conducted maintenance and studies

*The integrated luminosity for the period from 12/22/10 to 1/3/11 was ~53.9 inverse picobarns. NuMI reported receiving 8.8118 protons from 12/27/10 to 1/3/11.

Read the Current Accelerator Update
Read the Early Bird Report
View the Tevatron Luminosity Charts